40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Entry  Thu Aug 24 00:39:47 2017, gautam, Update, LSC, DRMI locking attempt 
    Reply  Fri Aug 25 01:20:52 2017, gautam, Update, LSC, DRMI locking attempt 
       Reply  Wed Aug 30 19:49:33 2017, gautam, Update, LSC, REFL55 demod board debugging REFL55_demod_check.pdf
          Reply  Thu Aug 31 00:52:52 2017, gautam, Update, LSC, REFL55 whitening board debugging REFL55_whtCheck.pdfREFL55_whtChk.tar.gz
             Reply  Thu Aug 31 03:31:15 2017, gautam, Update, LSC, DRMI re-locked! DRMI_1f.pngDRMI_relocked.pdf
                Reply  Mon Sep 4 16:30:06 2017, gautam, Update, LSC, Oplev loop tweaking 
                   Reply  Mon Sep 4 18:18:29 2017, rana, Update, LSC, dewhite switching: FOTON settings 
                   Reply  Tue Sep 5 02:07:49 2017, gautam, Update, LSC, Low Noise DRMI attempt MICH_err_comp.pdfdeWhitenedCoil.pdf
                      Reply  Wed Sep 6 23:06:30 2017, gautam, Update, LSC, Coil de-whitening switching investigation BScoils.pdfITMcoils.pdf
                         Reply  Fri Sep 15 16:00:33 2017, gautam, Update, LSC, Sensing measurement DRMI1f_Sep5.pdf
                         Reply  Fri Sep 15 17:08:58 2017, gautam, Update, LSC, Coil de-whitening switching investigation DRMI_IntensityNoise.pdf
                            Reply  Sat Sep 16 10:56:19 2017, rana, Update, LSC, Coil de-whitening switching investigation 
                               Reply  Fri Sep 22 18:12:27 2017, gautam, Update, LSC, DAC noise measurement (again) 
                                  Reply  Wed Sep 27 22:25:21 2017, gautam, Update, LSC, DAC noise measurement (again) coilNoises.pdf
                                     Reply  Thu Sep 28 06:35:44 2017, Chris, Update, LSC, DAC noise measurement (again) 
                                        Reply  Thu Sep 28 23:50:04 2017, gautam, Update, LSC, DAC noise measurement (again) 
                                           Reply  Fri Sep 29 18:36:25 2017, gautam, Update, LSC, DAC noise measurement (again) 
                               Reply  Tue Oct 3 01:32:39 2017, gautam, Update, LSC, Laser intensity noise coupling to MICH (simulated) MICH_intensityNoiseCoupling.pdfMICH_intensityNoiseASD.pdf
Message ID: 13338     Entry time: Thu Sep 28 06:35:44 2017     In reply to: 13336     Reply to this: 13343
Author: Chris 
Type: Update 
Category: LSC 
Subject: DAC noise measurement (again) 

Since you're monitoring two channels simultaneously, you could try subtracting them, as an alternative to carving out bandstops.

  1. Drive both channels with the same time series
  2. Tweak the filter module gains if needed to equalize the analog outputs
  3. Apply SR785 user math to subtract the two channels (or A-B mode of a single input, if you're not using that already?)
  4. Measure the residual, which should be the incoherent part containing DAC + coil driver noise

Subtraction can conceal certain annoying effects (like numerical noise or level crossing glitches) that remain coherent for two identical outputs. It might be worth experimenting with a differential offset or sinusoid, to try to break up that kind of coherence if it exists.

ELOG V3.1.3-